OpenStudy (anonymous):

what is the simplified form of :

OpenStudy (anonymous):

\[\frac{ 9x^2 - 25 }{ 3x+5 }\]

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

numerator: (3x)^2 -5^2 = (3x+5)(3x-5) cancel out with denominator

OpenStudy (kagıtucak):

\[(3x)^2-5^2/3x+5=(3x-5)(3x+5)/3x+5=\]

OpenStudy (kagıtucak):

3x-5

OpenStudy (anonymous):

thanks man